Help!! Someone I am having trouble with word problem, by turning this into an algebraic equation in order to solve it:
A university bookstore recently sold a wirebound graph-paper notebook for $2.50 and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of the notebooks were sold for a total of 118.60. How many of each type were sold? Thank you in advance.

Answer by oscargut(2103)

let x= number of graph-paper notebook for $2.50
let y= number of college-ruled notebook for $2.30
x+y=50
2.5x+2.3y=118.6
y=50-x
2.5x+2.3(50-x)=118.6
0.2x+115=118.6
0.2x=3.6
x=3.6/0.2=18
y=50-x=32

Answer: there are 18 graph-paper notebooks and 32 college-ruled notebooks